OpenStudy (anonymous):

Triangle ABC is similar to triangle DEF. What is EF? AB= 12 BC= 15 AC= 19, DE= 19.2 DF= 30.4

OpenStudy (anonymous):

can you show a figure please?

Directrix (directrix):

Corresponding sides of similar triangles are in proportion by definition of similar triangles. AB is to DE as BC is to EF AB/DE = BC/EF 12/19.2 = 15/EF 12* EF = 15 * 19.2 Solve for EF. Note: We knew which sides are corresponding because the correspondence of vertices was given in the similarity statement: Triangle ABC is similar to triangle DEF.

OpenStudy (anonymous):

Thanks to you mate know I understand. EF quals 24

Directrix (directrix):

That is what I got.